OpenStudy (anonymous):

f(x) = 5x − 7 g(x) = 3 − x^2 find (g o f)(x)

OpenStudy (anonymous):

how is it possible to substitute x for f(x)?

OpenStudy (mertsj):

\[g(f(x))=3-(5x-7)^2=3-(25x^2-70x+49)=-25x^2+70x-46\]

OpenStudy (mertsj):

g composite f means replace every x in the function g with the f(x)

OpenStudy (anonymous):

ohh, okay i understand now. i plugging the original equation into f(x)

OpenStudy (mertsj):

Good for you.

OpenStudy (anonymous):

i dont know if you are still here.. but where did the 70x come from?

OpenStudy (mertsj):

square 5x-7

OpenStudy (anonymous):

oh okay i need to do (5x-7)(5x-7)

OpenStudy (mertsj):